OpenStudy (anonymous):

Use a double angle formula to show that cot(pi/12) satisfies the equation c^2 -2sqrt3c -1 =0 Deduce that cot(pi/12) = 2+sqrt3

OpenStudy (john_es):

\[\tan\left(\frac{\pi}{12}\right)=\tan\left(\frac{1}{2}\frac{\pi}{6}\right)\] \[\tan\left(\frac{\pi}{12}+\frac{\pi}{12}\right)=(2\tan(\pi/12))/(1-\tan^2(\pi/12))\] \[\tan\left(\frac{\pi}{12}+\frac{\pi}{12}\right)=\tan(\pi/6)=1/\sqrt{3}\] Let \[c=1/\tan(\pi/12)\]Then \[1/\sqrt{3}=(2/c)/(1-1/c^2)\] \[c^2-2\sqrt{3}c-1=0\] Solving the quadratic equation (choosing the plus sign, to stay in the first quadrant) \[c=2+\sqrt{3}\]

OpenStudy (anonymous):

Brilliant! Thank you so much